OpenStudy (anonymous):

The equation of a circle is shown below. (x -5)2 + (y + 2)2 = 64 The radius of the circle is ________ units

OpenStudy (anonymous):

@Mousam

jimthompson5910 (jim_thompson5910):

hint: (x-h)^2 + (y-k)^2 = r^2 is a circle with a radius of r

jimthompson5910 (jim_thompson5910):

it also has the center of (h,k)

OpenStudy (anonymous):

ohh

jimthompson5910 (jim_thompson5910):

so what is r in this case

OpenStudy (anonymous):

sorry, i was helping someone else...

OpenStudy (anonymous):

64

jimthompson5910 (jim_thompson5910):

that's r^2

jimthompson5910 (jim_thompson5910):

if r^2 = 64, then what is r?

OpenStudy (anonymous):

64*64

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

no, you need to find a whole number such that when you multiply it by itself, you get 64

OpenStudy (anonymous):

@Scroll0405 since @jim_thompson5910 said that r^2 =64...to find r, we need to square root both side

jimthompson5910 (jim_thompson5910):

ex: if r^2 = 36 then r = 6 since r^2 = 6^2 = 6*6 = 36

OpenStudy (anonymous):

8*8

OpenStudy (anonymous):

=64

jimthompson5910 (jim_thompson5910):

good, so r = 8

OpenStudy (anonymous):

yup...

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

so is that the anwer

OpenStudy (anonymous):

yup the radius is 8